How fast (in rpm) must a centrifuge rotate if a particle 9.00 cm from the axis of rotation is to experience an acceleration of 121000 g's? If the answer has 4 digits or more, enter it without commas, e.g. 13500.
Answer #1

Centrifugal acceleration is given by the formula

a = w2 × r

r = 9 cm = 0.09m

a = 121000 × 9.8 m/s2= 1185800 m/s2

Hence, w = 1147 rad/s = 10961 rpm